mm Hg) obtained from the same woman. Find the regression equation, letting the right arm blood pressure be the predictor (x)Listed below are systolic blood pressure measurements​ (in mm​ Hg) obtained from the same woman. Find the regression​ equation, letting the right arm blood pressure be the predictor​ (x) variable. Find the best predicted systolic blood pressure in the left arm given that the systolic blood pressure in the right arm is 85 mm Hg. Use a significance level of 0.05.
Answer #1

Solution:

Here, we have to find the regression model for the prediction of the dependent variable or response variable systolic blood pressure in the left arm based on the independent variable or explanatory variable systolic blood pressure in the right arm.

The required regression model is given as below:

Regression Statistics

Multiple R

0.864491266

R Square

0.74734515

Adjusted R Square

0.663126866

Standard Error

9.084817982

Observations

5

ANOVA

df

SS

MS

F

Significance F

Regression

1

732.3982467

732.3982467

8.873906226

0.058648137

Residual

3

247.6017533

82.53391776

Total

4

980

Coefficients

Standard Error

t Stat

P-value

Lower 95%

Upper 95%

Intercept

68.98747652

31.81981253

2.168066718

0.118671555

-32.27736829

170.2523213

Right arm

1.070757671

0.359446101

2.978910241

0.058648137

-0.073160244

2.214675586

From above regression model, the p-value is given as 0.058648137 which is greater than α = 0.05, so we do not reject the null hypothesis. There is insufficient evidence to conclude that there is statistically significant relationship between the given two variables systolic blood pressure in the left arm and systolic blood pressure in the right arm.

The required regression equation is given as below:

ŷ = 69.0 + 1.1*x

Where, ŷ = Predicted systolic blood pressure in the left arm and x = systolic blood pressure in the right arm.

Now, we have to find the predicted systolic blood pressure in the left arm for the given systolic blood pressure of 85 mm Hg in the right arm.

So, we are given x = 85

ŷ = 69.0 + 1.1*x

ŷ = 69.0 + 1.1*85

ŷ = 162.5

Given that the systolic blood pressure in the right arm is 85 mm Hg, the best predicted systolic blood pressure in the left arm is 162.5 mm Hg.